# MNV-17: A High-Quality Performative Mandarin Dataset for Nonverbal Vocalization Recognition in Speech

### v1

@@ Abstract
Mainstream Automatic Speech Recognition (ASR) systems excel at transcribing lexical content, but largely fail to recognize nonverbal vocalizations (NVs) embedded in speech, such as sighs, laughs, and coughs. This capability is important for a comprehensive understanding of human communication, as NVs convey crucial emotional and intentional cues. Progress in NV-aware ASR has been hindered by the lack of high-quality, well-annotated datasets. To address this gap, we introduce MNV-17, a 7.55-hour performative Mandarin speech dataset. Unlike most existing corpora that rely on model-based detection, MNV-17’s performative nature ensures high-fidelity, clearly articulated NV instances. To the best of our knowledge, MNV-17 provides the most extensive set of nonverbal vocalization categories, comprising 17 distinct and well-balanced classes of common NVs. We benchmarked MNV-17 on four mainstream ASR architectures, evaluating their joint performance on semantic transcription and NV classification. The dataset and the pretrained model checkpoints will be made publicly available to facilitate future research in expressive ASR.
@@ 3.1 Joint ASR and NV Recognition
This experiment jointly evaluates both the ASR and NV recognition capabilities of the models. We finetuned four pretrained systems on the MNV-17 training set: SenseVoice [37], Paraformer (zh version), Qwen2-Audio (instruct version) [38], and Qwen2.5-Omini [39]. For each model, the checkpoint with the best performance on the validation set was selected for inference on the test set. Following standard practice [3, 40], punctuation and spaces were removed from the transcripts before calculating the Character Error Rate (CER). Each NV label was treated as a single character, and the final CER was averaged across all samples in the test set.
@@
Qwen2.5-Omini 3.60
The results reveal a performance difference that can be attributed to the models’ architectural designs. The autoregressive Large Audio Models, Qwen2-Audio and Qwen2.5-Omini, significantly outperform their non-autoregressive counterparts, SenseVoice and Paraformer. This advantage may arise from the inherent flexibility of the autoregressive framework. By generating the output sequence token by token, these models can more effectively integrate the recognition of non-lexical NV events alongside standard lexical transcription, as each new prediction is conditioned on the rich, mixed context of previously generated words and NV tags. In contrast, non-autoregressive models, which are optimized for efficient, parallelized prediction, might be less effective at capturing the complex, sequential interaction between semantic content and discrete nonverbal events within a unified output. The leading performance of Qwen2.5-Omini, which achieved the lowest CER of 3.60%, underscores the potential of large-scale autoregressive architectures for this complex joint task.
@@ 3.2 NV Recognition Accuracy
In contrast, the top-performing models, SenseVoice, Qwen2-Audio, and Qwen2.5-Omini, all were pretrained across diverse audio tasks. This foundational training equipped them with generalized acoustic representations and strong sensitivity to various acoustic events. Fine-tuning these robust models on MNV-17 effectively activated and refined their latent nonverbal recognition capabilities, demonstrating that our dataset excels at leveraging pre-existing multi-task knowledge for the specific challenge of expressive ASR.
@@ 3.3 Impact of NV Integration on ASR Performance
In contrast, models with multi-task pre-training backgrounds demonstrated far greater resilience. SenseVoice and Qwen2.5-Omini experienced only minor CER increases of 0.47% and 0.19%, respectively, indicating they can absorb the new capability with negligible impact. Most notably, the ASR performance of Qwen2-Audio actually improved, with its CER dropping from 3.05% to 2.60%.
@@ 4 Conclusion
In this paper, we addressed the gap in mainstream ASR systems’ ability to recognize nonverbal vocalizations (NVs) by introducing MNV-17, a 7.55-hour, high-quality performative Mandarin speech dataset. Featuring 17 distinct and well-balanced NV categories, MNV-17 provides a robust resource for developing more expressive ASR. Our experiments demonstrate that large, multi-task audio models are particularly adept at the joint task of NV recognition and speech transcription, with Qwen2.5-Omini achieving a superior joint CER of 3.60%. Notably, our findings indicate that integrating NV recognition capabilities does not degrade, and can even enhance, the core ASR performance of these advanced models. We will make the dataset and benchmarks publicly available to facilitate future research.

### v2

@@ Abstract
Mainstream Automatic Speech Recognition (ASR) systems excel at transcribing lexical content, but largely fail to recognize nonverbal vocalizations (NVs) embedded in speech, such as sighs, laughs, and coughs. This capability is important for a comprehensive understanding of human communication, as NVs convey crucial emotional and intentional cues. Progress in NV-aware ASR has been hindered by the lack of high-quality, well-annotated datasets. To address this gap, we introduce MNV-17, a 7.55-hour performative Mandarin speech dataset. Unlike most existing corpora that rely on model-based detection, MNV-17’s performative nature ensures high-fidelity, clearly articulated NV instances. To the best of our knowledge, MNV-17 provides the most extensive set of nonverbal vocalization categories, comprising 17 distinct and well-balanced classes of common NVs. We benchmarked MNV-17 on four mainstream ASR architectures, evaluating their joint performance on semantic transcription and NV classification. The dataset and pretrained model checkpoints are publicly available to facilitate future research in expressive ASR.
@@ 3.1 Joint ASR and NV Recognition
This experiment jointly evaluates both the ASR and NV recognition capabilities of the models. We finetuned four pretrained systems on the MNV-17 training set: SenseVoice [37], Paraformer (zh version), Qwen2-Audio (instruct version) [38], and Qwen2.5-omni [39]. For each model, the checkpoint with the best performance on the validation set was selected for inference on the test set. Following standard practice [3, 40], punctuation and spaces were removed from the transcripts before calculating the Character Error Rate (CER). Each NV label was treated as a single character, and the final CER was averaged across all samples in the test set.
@@
Qwen2.5-omni 3.60
The results reveal a performance difference that can be attributed to the models’ architectural designs. The autoregressive Large Audio Models, Qwen2-Audio and Qwen2.5-omni, significantly outperform their non-autoregressive counterparts, SenseVoice and Paraformer. This advantage may arise from the inherent flexibility of the autoregressive framework. By generating the output sequence token by token, these models can more effectively integrate the recognition of non-lexical NV events alongside standard lexical transcription, as each new prediction is conditioned on the rich, mixed context of previously generated words and NV tags. In contrast, non-autoregressive models, which are optimized for efficient, parallelized prediction, might be less effective at capturing the complex, sequential interaction between semantic content and discrete nonverbal events within a unified output. The leading performance of Qwen2.5-omni, which achieved the lowest CER of 3.60%, underscores the potential of large-scale autoregressive architectures for this complex joint task.
@@ 3.2 NV Recognition Accuracy
In contrast, the top-performing models, SenseVoice, Qwen2-Audio, and Qwen2.5-omni, all were pretrained across diverse audio tasks. This foundational training equipped them with generalized acoustic representations and strong sensitivity to various acoustic events. Fine-tuning these robust models on MNV-17 effectively activated and refined their latent nonverbal recognition capabilities, demonstrating that our dataset excels at leveraging pre-existing multi-task knowledge for the specific challenge of expressive ASR.
@@ 3.3 Impact of NV Integration on ASR Performance
In contrast, models with multi-task pre-training backgrounds demonstrated far greater resilience. SenseVoice and Qwen2.5-omni experienced only minor CER increases of 0.47% and 0.19%, respectively, indicating they can absorb the new capability with negligible impact. Most notably, the ASR performance of Qwen2-Audio actually improved, with its CER dropping from 3.05% to 2.60%.
@@ 4 Conclusion
In this paper, we addressed the gap in mainstream ASR systems’ ability to recognize nonverbal vocalizations (NVs) by introducing MNV-17, a 7.55-hour, high-quality performative Mandarin speech dataset. Featuring 17 distinct and well-balanced NV categories, MNV-17 provides a robust resource for developing more expressive ASR. Our experiments demonstrate that large, multi-task audio models are particularly adept at the joint task of NV recognition and speech transcription, with Qwen2.5-omni achieving a superior joint CER of 3.60%. Notably, our findings indicate that integrating NV recognition capabilities does not degrade, and can even enhance, the core ASR performance of these advanced models. The dataset and pretrained model checkpoints are publicly available to facilitate future research in expressive ASR.